# Towards Energy-Aware Federated Learning via MARL: A Dual-Selection Approach for Model and Client

### v1

@@ 5. Experimental Results
To evaluate the performance of our proposed method, we implemented the DR-FL algorithm using PyTorch (version 1.4.0). Similar to FedAvg, we assume that only 10% of AIoT devices were involved in each round of FL communication during the training period. For DR-FL and other heterogeneous FL methods, we set the small batch size to 32. The number of local training epochs and the initial learning rate were 5 and 0.05, respectively. To simulate a variety of energy-constrained scenarios, we assume that each device is powered by a battery with a maximum capacity of 7,560 joules. In other words, each battery capacity is 1500 mA at a rated voltage of 5.04V. We conducted comprehensive experiments to answer the following four Research Questions (RQs).
RQ1: (Superiority of DR-FL ): What advantages can DR-FL achieve compared with state-of-the-art heterogeneous FL methods?
RQ2: (Benefits of MARL-based Dual-Selection): What benefits does MARL-based Dual-Selection provide during DR-FL learning, especially under constraints such as device energy and overall training time, compared with other SOTA heterogenous FL methods?
RQ3: (Scalability of DR-FL): How does the number of AIoT devices participating in knowledge sharing affect the performance of DR-FL?
RQ4: (Exploration of the Validation Data Ratio): How does the proportion of validation data in MARL affect the performance of DR-FL?
@@ 5.1.1. Model Settings
We compared our DR-FL method with two typical state-of-the-art heterogeneous FL methods, i.e., HeteroFL (Diao et al., 2021) and ScaleFL (Ilhan et al., 2023), which belong to subnetwork aggregation-based methods and knowledge distillation-based methods, respectively. We set the ResNet-18 model (He et al., 2015) as the backbone, where each block of the ResNet-18 model is followed by a new pair of the bottleneck and classifier, thus forming four new heterogeneous layer-wise models to simulate four types of heterogeneous models (i.e., Models 1-4 shown in Table 1). Note that each layer-wise model can be reused with the same backbone for the purpose of model inference.
@@ 5.1.2. Dataset Settings
To evaluate the effectiveness of DR-FL, we considered four training datasets: i.e., CIFAR10, CIFAR100 (Krizhevsky, 2009), Street View House Numbers (SVHN) (Netzer et al., 2011), Fashion-MNIST (Xiao et al., 2017). CIFAR10: The CIFAR10 dataset consists of 60,000 $32\times 32$ colour images across ten classes, with 6,000 images per class. The dataset is split into 50,000 training images and 10,000 testing images. CIFAR100: The CIFAR100 dataset is similar to CIFAR10 but contains 100 classes instead of 10, with 600 images per class. The dataset also comprises 50,000 training images and 10,000 testing images. SVHN: The SVHN dataset is a real-world image dataset derived from house numbers in Google Street View images. It contains over 600,000 labelled digit images, where each image is a 32×32 colour image representing a single digit (0-9). Fashion-MNIST: The Fashion-MNIST dataset is a dataset of Zalando’s article images, consisting of 70,000 $28\times 28$ grayscale images of 10 different fashion categories. In subsequent experiments, we investigated three non-Independent and Identically Distributed (non-IID) distributions for each dataset. Similar to the work of HeteroFL in (Diao et al., 2021), we constructed non-IID local training datasets using heterogeneous data splits following a Dirichlet distribution controlled by a variable $\alpha$. Typically, a smaller value of $\alpha$ represents a higher degree of a corresponding non-IID distribution. Meanwhile, we used the same data augmentation technologies to fully utilize natural image datasets as the ones used in HeteroFL (Diao et al., 2021). To enable MARL training on a cloud server in DR-FL, we used 4% of the overall training data as the validation set on the server. Note that the validation set on the server does not overlap with local training datasets hosted by AIoT devices.
@@
5.1.3. Test-bed Settings
Besides simulation-based evaluation, we constructed a physical test-bed platform as shown in Figure 4 to check the performance of our DR-FL in a real-world environment. The test-bed consists of four parts: i) the cloud server that is built on top of an Ubuntu workstation equipped with an Intel i9 CPU, 32G memory, and a GTX3090 GPU; ii) the Jetson Nano boards, where each of them has a quad-core ARM A57 CPU, a 128-core NVIDIA Maxwell GPU, and 4GB LPDDR4 RAM; iii) the Jetson AGX Xavier boards, where each of them is equipped with an 8-core CPU and a 512-core Volta GPU; and iv) an HP 9800 power meter (see the top-left part in Figure 4(a)) produced by Shenzhen HOPI Electronic Technology Ltd. Note that, along with the federated training process, we used the power meter to record the energy consumption of all the AIoT devices every second for the MARL environment construction.
@@ 5.2. Accuracy Comparison (RQ1)
To evaluate the effectiveness of our proposed DR-FL, Table 1 presents the best test accuracy information for HeteroFL, ScaleFL and our DR-FL under the specific energy constraints along the FL processes based on the four datasets, assuming all the device batteries are initialized to be full. For each dataset and FL method combination, we considered three kinds of data distributions for all local AIoT devices, where the non-IID settings follow the Dirichlet distributions controlled by $\alpha$. Note that the baseline approaches (HeteroFL and ScaleFL) do not consider the energetic constraints in their FL procedure. To make a fair comparison, we added the greedy algorithm for energy awareness in this experiment (model selection will select the maximum model that can be trained for FL) into the two baseline algorithms for comparison. The experiments were repeated five times to calculate the mean and variance.
@@ 5.3. Comparison of Energy Consumption (RQ2)
To validate the performance of our DR-FL method in terms of energy consumption and running time, we conducted an experiment involving a total of 40 devices (i.e., 20 Jetson Nano boards and 20 AGX Xavier boards). Figure 5 compares the total remaining energy variation and running time in the federated training processes using HeteroFL (ScaleFL is the same energy consumption and running time in the greedy algorithm ) and DR-FL, respectively. For each subfigure, we use the notion $X\_Y$ to represent the total result of all the devices of type $Y$ using method $X$. If $Y$ is omitted, the notion $X$ denotes the total result involving all the devices. For example, in Figure 5(a), the legend DR-FL denotes the overall remaining energy of all 40 devices, while DR-FL_Nano represents the overall remaining energy of all the 20 Jetson Nano boards.
@@
Figure 5(a) shows that our method can have more training rounds under the same energy constraints, thus leading to better overall test accuracy and energy efficacy. For example, for HeteroFL, the Jetson AGX Xavier-based devices ran out of batteries in the 12th round. However, for DR-FL, the Jetson AGX Xavier-based devices ran out of batteries in the 18th round. Moreover, in Figure 5(b), we can clearly find an inflexion point in the 12th round for HeteroFL, after which only Jetson Nano-based devices are involved in federated training. However, for DR-FL, we can observe an inflexion point in the 15th round, indicating the effectiveness of the MARL algorithm in controlling the energy waste of the device with reduced useless wait and training time.
@@ 5.4. Scalability Analysis (RQ3)
Figure  6 compares the test accuracy of three methods (i.e., HeteroFL, ScaleFL, and DR-FL) for various non-IID scenarios with different numbers of devices under specific energy constraints. From this figure, we can observe that when more heterogeneous devices participate in FL, the superiority of DR-FL becomes more significant than that of the other two methods. For example, for the non-IID scenario of CIFAR10 (with $\alpha$=0.1), DR-FL consistently achieves higher test accuracy than ScaleFL and HeteroFL.
@@ 6. Conclusion
Federated Learning (FL) is expected to enable privacy-preserving collaborative learning among Artificial Intelligence of Things (AIoT) devices. However, due to various heterogeneous settings (e.g., non-IID data, device models with different architectures) and device resource constraints (e.g., computing power and energy capacity), existing FL-based AIoT design greatly suffers from the problems of low inference accuracy, rapid battery consumption and long training time. To address these issues, this paper introduces a novel FL framework that enables efficient knowledge sharing between heterogeneous devices under specific energy constraints. Based on our proposed layer-wise aggregation method and MARL-based dual selection mechanism, AIoT devices with different computational and energy capabilities can adaptively select appropriate local models to participate in global model training, where devices can effectively learn from each other through appropriate parts belonging to different layer-wise models. Comprehensive experiments performed on well-known datasets demonstrate the effectiveness of DR-FL for inference performance, energy consumption, and scalability.

### v2

@@ 5. Experimental Results
In order to evaluate the effectiveness of the approach we propose, we utilized the DR-FL algorithm by employing PyTorch with version 1.4.0. Like FedAvg, we make the assumption that only 10% of AIoT devices participated in each round of FL communication throughout the training period. Regarding the case of DR-FL and other heterogeneous FL algorithms, we assign a small batch size of 32. The local training epochs and initial learning rate were set at 5 and 0.05, respectively. In order to model a range of energy-limited situations, we assume that every gadget is equipped with a battery with a maximum capacity of 7,560 joules. To be precise, the capacity of each battery is 1500 mA with a rated voltage of 5.04V. We conducted extensive experiments to address the following four Research Questions (RQs).
RQ1: (Superiority of DR-FL): What advantages can DR-FL achieve compared to state-of-the-art heterogeneous FL methods?
RQ2: (Advantages of MARL-based Dual-Selection?) What advantages does MARL-based Dual-Selection offer in DR-FL procedure, particularly when dealing with limitations such as device energy and total training time, in comparison to other state-of-the-art heterogeneous FL methods?
RQ3: (Scalability of DR-FL): What is the impact of the quantity of AIoT devices engaged in knowledge sharing on the performance of DR-FL?
RQ4: (Investigation of the Validation Data Ratio): What is the impact of varying the proportion of validation data in MARL on the performance of DR-FL?
@@ 5.1.1. Model Settings
We conducted a comparison between our DR-FL approach and two well-known state-of-the-art heterogeneous FL methods, namely HeteroFL (Diao et al., 2021) and ScaleFL (Ilhan et al., 2023). HeteroFL falls under the category of subnetwork aggregation-based methods, while ScaleFL belongs to the knowledge distillation-based methods. The ResNet-18 model (He et al., 2015) serves as the backbone. Each block of the ResNet-18 model is accompanied by a bottleneck and classifier, resulting in the creation of four distinct layer-wise models. These models are designed to simulate four different types of heterogeneous models, referred to as Models 1-4 in Table 1. Note that each layer-wise model can be reused with the same backbone for the purpose of model inference.
@@ 5.1.2. Dataset Settings
To evaluate the effectiveness of DR-FL, we considered four training datasets: i.e., CIFAR10, CIFAR100 (Krizhevsky, 2009), Street View House Numbers (SVHN) (Netzer et al., 2011), Fashion-MNIST (Xiao et al., 2017). CIFAR10: The CIFAR10 dataset consists of 60,000 $32\times 32$ colour images across ten classes, with 6,000 images per class. The dataset is split into 50,000 training images and 10,000 testing images. CIFAR100: The CIFAR100 dataset is similar to CIFAR10 but contains 100 classes instead of 10, with 600 images per class. The dataset also comprises 50,000 training images and 10,000 testing images. SVHN: The SVHN dataset is a real-world image dataset derived from house numbers in Google Street View images. It contains over 600,000 labelled digit images, where each image is a 32×32 colour image representing a single digit (0-9). Fashion-MNIST: The Fashion-MNIST dataset is a dataset of digital number images (Xiao et al., 2017), consisting of 70,000 $28\times 28$ grayscale images of 10 different fashion categories. In our subsequent assessments, we evaluated three non-Independent and Identically Distributed (non-IID) distributions for each dataset. Following the methodology of HeteroFL as described in (Diao et al., 2021), we generated non-IID local training datasets by utilizing heterogeneous data splits. These splits were created based on a Dirichlet distribution, which was controlled by a variable $\alpha$. A smaller value of $\alpha$ often indicates a greater level of non-IID distribution. Additionally, we employed the same data augmentation techniques as those utilized in HeteroFL (Diao et al., 2021) to maximize the efficient use of natural picture datasets. To facilitate Multi-Agent Reinforcement Learning (MARL) training on a cloud server in DR-FL, we allocated 4% of the total training data as the validation set on the server. It is important to understand that the validation set used on the server is completely separate from the local training datasets stored on AIoT devices.
5.1.3. Test-bed Settings
In addition to conducting simulation-based evaluation, we developed a physical test-bed platform, depicted in Figure 4, to assess the effectiveness of our DR-FL in a real-world setting. The test-bed platform comprises four components: i) The cloud server is constructed on an Ubuntu workstation that has an Intel i9 CPU, 32G memory, and a GTX3090 GPU. ii) The Jetson Nano boards each contain a quad-core ARM A57 CPU, a 128-core NVIDIA Maxwell GPU, and 4GB LPDDR4 RAM. iii) The Jetson AGX Xavier boards are equipped with an 8-core CPU and a 512-core Volta GPU. iv) Shenzhen HOPI Electronic Technology Ltd. produces the HP 9800 power meter, which is located in the top-left part of Figure 4(a). It is important to mention that, in addition to the federated training process, we utilized a power meter to accurately measure the energy consumption of all the AIoT devices at one-second intervals during the development of the MARL environment.
@@ 5.2. Accuracy Comparison (RQ1)
To evaluate the effectiveness of our proposed DR-FL, Table 1 presents the best test accuracy information for HeteroFL, ScaleFL and our DR-FL under the specific energy constraints along the FL processes based on the four datasets, assuming all the device batteries are initialized to be full. For each dataset and FL method combination, we considered three kinds of data distributions for all local AIoT devices, where the non-IID settings follow the Dirichlet distributions controlled by $\alpha$. Note that the baseline approaches (HeteroFL and ScaleFL) do not consider the energetic constraints in their FL procedure. To make a fair comparison, we added the greedy algorithm for energy awareness in this experiment (model selection will select the maximum model that can be trained in FL) into the two baseline algorithms for comparison. The experiments were repeated five times to calculate the mean and variance.
@@ 5.3. Comparison of Energy Consumption (RQ2)
To evaluate the effectiveness of our DR-FL technique in terms of energy usage and execution time, we carried out an experiment comprising a total of 40 devices, specifically 20 Jetson Nano boards and 20 AGX Xavier boards. Figure 5 illustrates the differences in total remaining energy variation and running time between the federated learning processes employing HeteroFL (where ScaleFL has the same energy consumption and running time as the greedy algorithm) and DR-FL. Each subfigure is represented by the notation $X\_Y$, which denotes the cumulative outcome of all devices of type $Y$ employing technique $X$. If the variable $Y$ is excluded, the term $X$ represents the overall outcome that encompasses all the devices. For instance, in Figure 5(a), the label DR-FL indicates the total remaining energy of the 40 devices, whereas DR-FL_Nano represents the total remaining energy of the 20 Jetson Nano boards.
@@
Figure 5(a) demonstrates that our approach can accommodate a greater number of training rounds while adhering to the same energy limitations. Consequently, this results in improved test accuracy and energy efficiency. As an illustration, in the case of HeteroFL, the devices powered by Jetson AGX Xavier became depleted of battery power by the 12th round. Nevertheless, during the DR-FL event, the devices powered by Jetson AGX Xavier experienced battery depletion by the 18th round. Furthermore, in Figure 5(b), there is a distinct inflexion point observed in the 12th round for HeteroFL. Subsequently, only devices based on Jetson Nano are utilized for federated training. Nevertheless, in the case of DR-FL, a notable turning point can be observed in the 15th cycle, which signifies the efficacy of the MARL algorithm in managing the energy wastage of the device by minimizing useless waiting and training time.
@@ 5.4. Scalability Analysis (RQ3)
Figure  6 illustrates the test accuracy of three approaches (HeteroFL, ScaleFL, and DR-FL) in various non-IID scenarios with different numbers of devices, all within specified energy constraints. From this figure, we can observe that when more heterogeneous devices participate in FL, the superiority of DR-FL becomes more significant than that of the other two methods. For example, for the non-IID scenario of CIFAR10, Fashion-MNIST and SVHN (with $\alpha$=0.1), DR-FL consistently achieves higher test accuracy than ScaleFL and HeteroFL.
@@ 6. Conclusion
Federated Learning (FL) is intended to facilitate privacy-preserving collaborative learning among Artificial Intelligence of Things (AIoT) devices. Nevertheless, the current design of AIoT systems based on Federated Learning (FL) encounters significant challenges, such as non-IID data, heterogeneous local devices, and varying computational and energy capabilities. Consequently, these challenges result in issues such as low inference accuracy, excessive battery consumption, and increased training time. This paper presents an innovative Federated Learning (FL) paradigm that facilitates effective information exchange among various devices while considering unique energy limitations. Our proposed layer-wise aggregation method and MARL-based dual selection mechanism enable AIoT devices with varying computational and energy capabilities to intelligently choose suitable local models for global model training. This allows devices to effectively learn from each other by utilizing relevant components from different layer-wise models. The efficacy of DR-FL in terms of inference performance, energy consumption, and scalability has been demonstrated through extensive experiments conducted on widely recognized datasets.